when you double a number & add 22, you get 5/8 of the original number-what is the original number?
Answer by malakumar_kos@yahoo.com(315)


when you double a number & add 22, you get 5/8 of the original number-what is the original number
Let the original number be = x
When it is doubled it is = 2x
5/8 of the number is = 5x/8
Therefore eq'n is 2x+22 = 5x/8
8(2x+22) = 5x
16x+176 = 5x
16x-5x = -176
11x = -176
x = -176/11 = -16
The original number is -16